5(2x-1)-2x=11+2(3x-3)
Answer by fractalier(6550)

To solve this
5(2x-1)-2x=11+2(3x-3)
let us distribute and collect like terms...
10x - 15 - 2x = 11 + 6x - 6
8x - 15 = 6x + 5
Now subtract 6x and add 15 to both sides and get
2x - 15 = 5
2x = 20
x = 10